# Distributionally Robust Optimization with Adversarial Data Contamination

### v1

@@ A Bounded Geometric Search with Early Stopping
The geometric search ensures that for at least one candidate $D_{j}$, we have $D_{0}\leq D_{j}<2D_{0}$. For this $D_{j}$, running Algorithm 1 with $\gamma_{j}=D_{j}/(\zeta\sqrt{N})$ yields a solution $\hat{{\bm{w}}}_{T}^{(j)}$. As shown in the analysis of the error term in Proposition 4.1 when $\gamma$ is mismatched (see Equation 15), the function value suboptimality $O(D_{0}\Delta)$ becomes $O((D_{j}+D_{0}^{2}/D_{j})\Delta)$. Since $D_{j}\in[D_{0},2D_{0}]$, this is again $O(D_{0}\Delta)=O(\|{\bm{w}}_{0}-{\bm{w}}^{*}\|_{2}\sigma\zeta\sqrt{\epsilon})$.
@@
If the true $D_{0}$ is even smaller than $\Delta/\zeta$, our search does not try $D_{j}$ values below $\Delta/\zeta$. For the first run $D_{\min}=\Delta/\zeta$, the proven error bound (see Equation 15) depends on $(D_{\min}+D_{0}^{2}/D_{\min})\Delta$ and would be approximately $(\Delta/\zeta+D_{0}^{2}\zeta/\Delta)\Delta=\Delta^{2}/\zeta+D_{0}^{2}\zeta$. Since $D_{0}<\Delta/\zeta$, $D_{0}^{2}\zeta<(\Delta^{2}/\zeta^{2})\zeta=\Delta^{2}/\zeta$. So the error in this case is $O(\Delta^{2}/\zeta)=O(\sigma^{2}\zeta\epsilon)$.
@@
The final excess risk bound for the selected solution $\hat{{\bm{w}}}_{T}^{(*)}$ remains
$O(D_{0}\Delta+E_{\text{eval}})=O(\zeta(\|{\bm{w}}_{0}\|_{2}+\|{\bm{w}}^{*}\|_{2})\sigma\sqrt{\epsilon}+\sigma^{2}\zeta\epsilon)\;.$
Initializing at ${\bm{w}}_{0}={\bm{0}}$, this simplifies to $O(\zeta(\|{\bm{w}}^{*}\|_{2}+\sigma\sqrt{\epsilon})\sigma\sqrt{\epsilon})$.

### v2

@@ 4.2.4 Overall Error Guarantee and Complexity
The geometric search ensures that for at least one candidate $D_{j}$, we have $D_{0}\leq D_{j}<2D_{0}$. For this $D_{j}$, running Algorithm˜1 with $\gamma_{j}=D_{j}/(\zeta\sqrt{N})$ yields a solution $\hat{{\bm{w}}}_{T}^{(j)}$. As shown in the analysis of the error term in Proposition˜4.1 when $\gamma$ is mismatched (see Equation˜12), the function value suboptimality $O(D_{0}\Delta)$ becomes $O((D_{j}+D_{0}^{2}/D_{j})\Delta)$. Since $D_{j}\in[D_{0},2D_{0}]$, this is again $O(D_{0}\Delta)=O(\|{\bm{w}}_{0}-{\bm{w}}^{*}\|_{2}\sigma\zeta\sqrt{\epsilon})$.
@@
If the true $D_{0}$ is even smaller than $\Delta/\zeta$, our search does not try $D_{j}$ values below $\Delta/\zeta$. For the first run $D_{\min}=\Delta/\zeta$, the proven error bound (see Equation˜12) depends on $(D_{\min}+D_{0}^{2}/D_{\min})\Delta$ and would be approximately $(\Delta/\zeta+D_{0}^{2}\zeta/\Delta)\Delta=\Delta^{2}/\zeta+D_{0}^{2}\zeta$. Since $D_{0}<\Delta/\zeta$, $D_{0}^{2}\zeta<(\Delta^{2}/\zeta^{2})\zeta=\Delta^{2}/\zeta$. So the error in this case is $O(\Delta^{2}/\zeta)=O(\sigma^{2}\zeta\epsilon)$.
@@
The final excess risk bound for the selected solution $\hat{{\bm{w}}}_{T}^{(*)}$ remains $O(D_{0}\Delta+E_{\text{eval}})=O(\zeta(\|{\bm{w}}_{0}\|_{2}+\|{\bm{w}}^{*}\|_{2})\sigma\sqrt{\epsilon}+\sigma^{2}\zeta\epsilon)$. Initializing at ${\bm{w}}_{0}={\bm{0}}$, this simplifies to $O(\zeta(\|{\bm{w}}^{*}\|_{2}+\sigma\sqrt{\epsilon})\sigma\sqrt{\epsilon})$.